OpenStudy (anonymous):

If (3+2i)+(2+bi) = 5-4i, the value of b is a. -6 b. -2 c. 2 d. 6

OpenStudy (anonymous):

B = -6

Parth (parthkohli):

Hi! Can you simplify the left hand side?

OpenStudy (anonymous):

For LHS, combine like terms to get: 3 + 2 + 2i + bi = 5 + (2 + b)i Then, set the corresponding coefficients with i's: 2 + b = -4 Therefore... b = -4 - 2 b = -6

Parth (parthkohli):

Adding the reals, we get 5. Adding the imaginaries, we get (2 + b)i.

OpenStudy (anonymous):

then??? what after getting that???